If some neon gas at 121,000 Pa were allowed to expand from 37 mL to 60 mL without changing the temperature, what pressure would the neon gas exert under these new conditions?
1
Expert's answer
2020-05-15T10:22:48-0400

pV=nRT.

n=0.121 x 37 / 8.314 x 273. = 0,00197248826.

p= nRT/V=0,00197248826 x 8.314 x 273 / 60 = 0,0746166667 kpa.

74.616 pa.